# Optimal Denial-of-Service Attacks Against Partially-Observable Real-Time Monitoring Systems

### v1

@@ II-E Closed-Form Expression of the Optimal Solution
By definition of $\lambda(s_{k})$, $\overline{s}_{k}-\lambda(s_{k})\overline{a}_{k}=\overline{s}_{k+1}-\lambda(s_{k})\overline{a}_{k+1}$. Therefore, we have that:
$(\overline{s}_{k}-\lambda\overline{a}_{k})-(\overline{s}_{k+1}-\lambda\overline{a}_{k+1})$
$\qquad=\lambda(s_{k})\overline{a}_{k}-\lambda(s_{k})\overline{a}_{k+1}-\lambda\overline{a}_{k}+\lambda\overline{a}_{k+1}$
$\qquad=(\lambda-\lambda(s_{k}))(\overline{a}_{k+1}-\overline{a}_{k}).$
Given that $\overline{a}_{k}$ is strictly decreasing with respect to $k$, we can write $\overline{s}_{k}-\lambda\overline{a}_{k}\geq\overline{s}_{k+1}-\lambda\overline{a}_{k+1}$ when $\lambda\leq\lambda(s_{k})$, and $\overline{s}_{k}-\lambda\overline{a}_{k}<\overline{s}_{k+1}-\lambda\overline{a}_{k+1}$ when $\lambda>\lambda(s_{k})$. This completes the proof of the first and second conditions.
In order to prove the third condition, it is sufficient to show that $\overline{s}_{k}-\lambda\overline{a}_{k}$ is strictly increasing with respect to $k$ when $k\leq n$ and $\lambda>\lambda(s_{n})$. To that end, we prove that $\overline{s}_{k+1}-\lambda\overline{a}_{k+1}>\overline{s}_{k}-\lambda\overline{a}_{k}$ when $k\leq n$ and $\lambda>\lambda(s_{n})$. As $\lambda(s_{n})\geq\lambda(s_{k})$ by Lemma 3, we get $\lambda>\lambda(s_{k})$. Hence, according to (30), $\overline{s}_{k}-\lambda\overline{a}_{k}<\overline{s}_{k+1}-\lambda\overline{a}_{k+1}$. Thus, $\overline{s}_{k}-\lambda\overline{a}_{k}$ is strictly increasing with respect to $k$ when $k\leq n$ and $\lambda>\lambda(s_{n})$, which implies that $\overline{s}_{n+1}-\lambda\overline{a}_{n+1}>\overline{s}_{k}-\lambda\overline{a}_{k}$ for all $k\leq n$.
Finally, to prove the last condition, we show that $\overline{s}_{k}-\lambda\overline{a}_{k}$ is decreasing with respect to $k$ when $k>n$ and $\lambda\leq\lambda(s_{n+1})$. As $\lambda(s_{n+1})\leq\lambda(s_{k})$ by Lemma 3, we get $\lambda\leq\lambda(s_{k})$. Hence, according to (29), $\overline{s}_{k}-\lambda\overline{a}_{k}\geq\overline{s}_{k+1}-\lambda\overline{a}_{k+1}$. Thus, $\overline{s}_{k}-\lambda\overline{a}_{k}$ is decreasing with respect to $k$ when $k>n$ and $\lambda\leq\lambda(s_{n+1})$, which implies that $\overline{s}_{n+1}-\lambda\overline{a}_{n+1}\geq\overline{s}_{k}-\lambda\overline{a}_{k}$ for all $k>n$. ∎
@@
When $\lambda\leq\lambda(s_{0})$, we have $\lambda\leq\lambda(s_{k})$ for all $k\geq 0$. Hence, from (29), we get $\overline{s}_{k}-\lambda\overline{a}^{k}\geq\overline{s}_{k+1}-\lambda\overline{a}_{k+1}$. Hence, in this regime for $\lambda$, we can deduce that the optimal threshold value is $s_{0}$, as $\overline{s}_{0}-\lambda\overline{a}_{0}\geq\overline{s}_{k+1}-\lambda\overline{a}_{k+1}$. That completes the proof of the first statement.
However, when $\lambda(n)<\lambda\leq\lambda(n+1)$, from (31) and (32), we obtain that $\overline{s}_{n+1}-\lambda\overline{a}_{n+1}\geq\overline{s}_{k}-\lambda\overline{a}_{k}$, for all $k\geq 0$. Therefore, in this regime for $\lambda$, we can deduce that the optimal threshold value is $s_{n+1}$. That completes the proof of the second statement.
Let $\lim_{k\to\infty}\lambda(s_{k})=\lambda_{\infty}$. Since $\lambda(.)$ is increasing in $s_{k}$, then $\lambda(s_{k})<\lambda_{\infty}$ for all $k$. Thus, when $\lambda_{\infty}\leq\lambda$, from (30), we have $\overline{s}_{k}-\lambda\overline{a}_{k}<\overline{s}_{k+1}-\lambda\overline{a}_{k+1}$ as $\lambda(s_{k})<\lambda_{\infty}\leq\lambda$ for all $k$. Note that the optimal threshold $s_{n}$ cannot be finite, otherwise $\overline{s}_{n}-\lambda\overline{a}_{n}\geq\overline{s}_{n+1}-\lambda\overline{a}_{n+1}$, which is a contradiction. That completes the proof of the third statement. ∎
@@
Based on Theorem 2, we provide Algorithm 1 for the calculation of the optimal threshold value.
Algorithm 1 Optimal Jamming Policy
1:Input $\lambda$ and $\lambda(\cdot)$
2:Initialization $t=0$, $x_{0}=0$, and $k=1$
3:if $\lambda\leq\lambda(s_{0})$ then: $n^{*}=0$
4:else if $\lambda(s_{n})\geq\lim_{k\rightarrow\infty}\lambda(s_{k})$ then: $n^{*}=\infty$
5:else
6:     while $k==1$ do $x_{t+1}=x_{t}+\alpha(\lambda-\lambda(s_{x_{t}}))$
7:         if $\lambda(s_{\lfloor x_{t+1}\rfloor})<\lambda\leq\lambda(s_{\lfloor x_{t+1}\rfloor+1})$ then
8:              $k=0$ and $n^{*}=\lfloor x_{t+1}\rfloor+1$
9:         end if
10:     end while
11:end if
12:return $s_{n^{*}}$
Note that, in Algorithm 1, $\lfloor x\rfloor$ represents the integer part of $x$, $\alpha$ is given in [26, Theorem 7], and the continuous extension of the function $\lambda(.)$ in $\mathbb{R}^{+}$ is based on the linear interpolation
$\lambda(s_{x})=\left\{\begin{array}[]{ll}\lambda(s_{i})&\text{if}\ x=i\in\mathbb{N},\\[4.5pt] \lambda(s_{i+1})(x-i)\\[4.5pt] \quad-\lambda(s_{i})(x-i-1)&\text{if}\ x\in[i,i+1].\end{array}\right.$ (33)
@@ III-A Problem Statement
Let $N$ be the total number of subsystems, $p_{i}$ be the probability that the packet sent by the source $i$ is successfully decoded, $r_{i}$ be the probability that the source $i$ transitions from one state to another, and $q_{i}$ be the probability that the adversary succeeds in jamming the channel $i$. Due to the energy constraint, the adversary can jam only $M<N$ channels. A jamming policy $\phi$ is defined as a sequence of actions $\phi=(\boldsymbol{a}_{\phi}(0),\boldsymbol{a}_{\phi}(1),\ldots)$, where $\boldsymbol{a}_{\phi}(t)=(a^{1}_{\phi}(t),a^{2}_{\phi}(t),\ldots,a^{N}_{\phi}(t))$ is a binary vector such that $a^{i}_{\phi}(t)=1$ if channel $i$ is jammed at time $t$. Let $\boldsymbol{s}(t)=(s^{1}(t),\ldots,s^{N}(t))$ be a vector at time $t$ such that $s^{i}(t)$ is the EAoII of subsystem $i$ at time $t$. Our aim is to find the optimal jamming policy that maximizes the total average EAoII subject to the constraint on the number of channels under attack. This can be formulated as
@@
$\text{subject to}\ \sum_{i=1}^{N}a^{i}_{\phi}(t)\leq\alpha N\quad\forall t\geq 0$ (34b)
@@ III-B Lagrangian Relaxation
subject $\text{subject to}\ \lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{\phi}\Big{[}\sum_{t=0}^{T-1}\sum_{i=1}^{N}a^{i}_{\phi}(t)\Big{]}\leq\alpha N.$ (35b)
@@
$f(W,\phi)$ $=\!\lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{\phi}\Big{[}\sum_{t=0}^{T-1}\sum_{i=1}^{N}s^{i}_{\phi}(t)\!-\!Wa^{i}_{\phi}(t)\big{|}\boldsymbol{s}(0)\Big{]}$
@@
$\underset{\phi\in\Phi}{\text{maximize}}\ \lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{\phi}\Big{[}\sum_{t=0}^{T-1}s^{i}_{\phi}(t)-Wa^{i}_{\phi}(t)\big{|}s^{i}(0)\Big{]}.$ (38)
@@ III-C Whittle’s Index Jamming Policy
and $a_{n}^{i}(W)$ be the action at state $s_{n}^{i}$ associated with subsystem $i$ under the optimal threshold jamming policy given the Lagrangian parameter $W$. We first consider the steady-state form of the problem in (38) under the threshold jamming policy with parameter $s^{i}_{n}$, i.e.,
$\underset{n\in\mathbb{N}}{\text{maximize}}\ \overline{s}_{n}^{i}-W\overline{a}_{n}^{i}$ (40)
where $\overline{s}_{n}^{i}$ and $\overline{a}_{n}^{i}$ are given in (2) and (26), respectively, when replacing $p$, $q$, and $r$ with $p_{i}$, $q_{i}$, $r_{i}$. To ensure the existence of the Whittle’s indices, it is imperative to establish the indexability property for all subsystems. To accomplish this, we initially formalize the concepts of indexability and the Whittle’s index in the following definitions.
@@
In the context of the problem in (38), for a given $W$, we define $D^{i}(W)$ as the set of states in which the optimal action for each subsystem $i$ is passive, i.e., $a^{i}_{n}(W)=0$. In other words, $s^{i}_{n}\in D^{i}(W)$ if and only if the optimal action at state $s^{i}_{n}$ is passive.
@@
To prove the claim, it is sufficient to show that $\overline{a}^{i}_{n}$ is decreasing with respect to $n$. Indeed, we can write
$\overline{a}^{i}_{n+1}-\overline{a}^{i}_{n}$
@@
Now that the indexability property has been established, we can assert the existence of the Whittle’s index. Building on the procedure introduced in [29, Algorithm 1], we propose a tailored algorithm that iteratively computes the Whittle’s index values for each subsystem $i$. The following lemma shows that the output of this algorithm is compatible with Definition 3.
@@
2:Find $W_{0}^{i}=\underset{n\in\mathbb{N}}{\text{inf}}\frac{\overline{s}_{n}^{i}-\overline{s}_{0}^{i}}{\overline{a}_{n}^{i}-\overline{a}_{0}^{i}}$
@@
8:     Find $W_{j}^{i}=\underset{n\in\mathbb{N}\setminus M_{j}^{i}}{\text{inf}}\frac{\overline{s}_{n}^{i}-\overline{s}_{n_{j}}^{i}}{\overline{a}_{n}^{i}-\overline{a}_{n_{j}}^{i}}$
@@
The proof follows directly the results in [29, Proposition 4]. ∎
In the next lemma, we provide a useful property that is satisfied by the sequences $\overline{s}_{k}^{i}$ and $\overline{a}_{k}^{i}$.
@@
For each subsystem $i$, if $\overline{a}_{k}^{i}$ is decreasing with $k$ and
$\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i}}<\frac{\overline{s}_{k+2}^{i}-\overline{s}_{k+1}^{i}}{\overline{a}_{k+2}^{i}-\overline{a}_{k+1}^{i}}$ (43)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{n}^{i}-\overline{a}_{k}^{i}}\geq\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i}}$ (44)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{n}^{i}-\overline{a}_{k}^{i}}=\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i}}\geq\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i}}$ (45)
@@
$\frac{\overline{s}_{n+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i}}=$ $\frac{\overline{s}_{n+1}^{i}-\overline{s}_{k}^{i}+\overline{s}_{n}^{i}-\overline{s}_{n}^{i}}{\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i}}$
$=$ $\frac{\overline{s}_{n+1}^{i}-\overline{s}_{n}^{i}}{\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i}}+\frac{\overline{s}_{n}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i}}$
$\geq$ $\underbrace{\frac{(\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i})(\overline{a}_{n+1}^{i}-\overline{a}_{n}^{i})}{(\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i})(\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i})}}_{(a)}+\underbrace{\frac{(\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i})(\overline{a}_{n}^{i}-\overline{a}_{k}^{i})}{(\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i})(\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i})}}_{(b)}$
$=$ $\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i}}\Big{[}\frac{\overline{a}_{n+1}^{i}-\overline{a}_{n}^{i}}{\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i}}+\frac{\overline{a}_{n}^{i}-\overline{a}_{k}^{i}}{\overline{a}_{n+1}^{i}-\overline{a}_{k}^{i}}\Big{]}$
$=$ $\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i}}$ (47)
where $(a)$ comes from (43) and the fact that $\overline{a}_{k}^{i}$ is strictly decreasing with respect to $k$, and (b) comes from (45) and the fact that $\overline{a}_{k}^{i}$ is decreasing with respect to $k$. Therefore, the inequality also holds for $n+1$. This concludes the proof. ∎
@@
$W(s_{n}^{i})=\lambda_{i}(s_{n})=\frac{\overline{s}_{n+1}^{i}-\overline{s}_{n}^{i}}{\overline{a}_{n+1}^{i}-\overline{a}_{n}^{i}}$ (48)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{j}^{i}}{\overline{a}_{n}^{i}-\overline{a}_{j}^{i}}\geq\frac{\overline{s}_{j+1}^{i}-\overline{s}_{j}^{i}}{\overline{a}_{j+1}^{i}-\overline{a}_{j}^{i}}$ (50)
@@
$\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{a}_{k+1}^{i}-\overline{a}_{k}^{i}}<\frac{\overline{s}_{k+2}^{i}-\overline{s}_{k+1}^{i}}{\overline{a}_{k+2}^{i}-\overline{a}_{k+1}^{i}}$ (51)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{j}^{i}}{\overline{a}_{n}^{i}-\overline{a}_{j}^{i}}\geq\frac{\overline{s}_{j+1}^{i}-\overline{s}_{j}^{i}}{\overline{a}_{j+1}^{i}-\overline{a}_{j}^{i}}$ (52)
for all $n>j$. Thus, the minimizer of $(\overline{s}_{n}^{i}-\overline{s}_{j}^{i})/(\overline{a}_{n}^{i}-\overline{a}_{j}^{i})$ over $n$ at step $j$ is $j+1$. As a result, the Whittle’s index at state $s_{j}^{i}$ is obtained as
$W(s_{j}^{i})=W_{j}^{i}=\frac{\overline{s}_{j+1}^{i}-\overline{s}_{j}^{i}}{\overline{a}_{j+1}^{i}-\overline{a}_{j}^{i}}=\lambda_{i}(s_{j}).$ (53)

### v2

@@ II-E Closed-Form Expression of the Optimal Solution
By definition of $\lambda(s_{k})$, $\overline{s}_{k}-\lambda(s_{k})\overline{d}_{k}=\overline{s}_{k+1}-\lambda(s_{k})\overline{d}_{k+1}$. Therefore, we have that:
$(\overline{s}_{k}-\lambda\overline{d}_{k})-(\overline{s}_{k+1}-\lambda\overline{d}_{k+1})$
$\qquad=\lambda(s_{k})\overline{d}_{k}-\lambda(s_{k})\overline{d}_{k+1}-\lambda\overline{d}_{k}+\lambda\overline{d}_{k+1}$
$\qquad=(\lambda-\lambda(s_{k}))(\overline{d}_{k+1}-\overline{d}_{k}).$
Given that $\overline{d}_{k}$ is strictly decreasing with respect to $k$, we can write $\overline{s}_{k}-\lambda\overline{d}_{k}\geq\overline{s}_{k+1}-\lambda\overline{d}_{k+1}$ when $\lambda\leq\lambda(s_{k})$, and $\overline{s}_{k}-\lambda\overline{d}_{k}<\overline{s}_{k+1}-\lambda\overline{d}_{k+1}$ when $\lambda>\lambda(s_{k})$. This completes the proof of the first and second conditions.
In order to prove the third condition, it is sufficient to show that $\overline{s}_{k}-\lambda\overline{d}_{k}$ is strictly increasing with respect to $k$ when $k\leq n$ and $\lambda>\lambda(s_{n})$. To that end, we prove that $\overline{s}_{k+1}-\lambda\overline{d}_{k+1}>\overline{s}_{k}-\lambda\overline{d}_{k}$ when $k\leq n$ and $\lambda>\lambda(s_{n})$. As $\lambda(s_{n})\geq\lambda(s_{k})$ by Lemma 3, we get $\lambda>\lambda(s_{k})$. Hence, according to (30), $\overline{s}_{k}-\lambda\overline{d}_{k}<\overline{s}_{k+1}-\lambda\overline{d}_{k+1}$. Thus, $\overline{s}_{k}-\lambda\overline{d}_{k}$ is strictly increasing with respect to $k$ when $k\leq n$ and $\lambda>\lambda(s_{n})$, which implies that $\overline{s}_{n+1}-\lambda\overline{d}_{n+1}>\overline{s}_{k}-\lambda\overline{d}_{k}$ for all $k\leq n$.
Finally, to prove the last condition, we show that $\overline{s}_{k}-\lambda\overline{d}_{k}$ is decreasing with respect to $k$ when $k>n$ and $\lambda\leq\lambda(s_{n+1})$. As $\lambda(s_{n+1})\leq\lambda(s_{k})$ by Lemma 3, we get $\lambda\leq\lambda(s_{k})$. Hence, according to (29), $\overline{s}_{k}-\lambda\overline{d}_{k}\geq\overline{s}_{k+1}-\lambda\overline{d}_{k+1}$. Thus, $\overline{s}_{k}-\lambda\overline{d}_{k}$ is decreasing with respect to $k$ when $k>n$ and $\lambda\leq\lambda(s_{n+1})$, which implies that $\overline{s}_{n+1}-\lambda\overline{d}_{n+1}\geq\overline{s}_{k}-\lambda\overline{d}_{k}$ for all $k>n$. ∎
@@
When $\lambda\leq\lambda(s_{0})$, we have $\lambda\leq\lambda(s_{k})$ for all $k\geq 0$. Hence, from (29), we get $\overline{s}_{k}-\lambda\overline{d}^{k}\geq\overline{s}_{k+1}-\lambda\overline{d}_{k+1}$. Hence, in this regime for $\lambda$, we can deduce that the optimal threshold value is $s_{0}$, as $\overline{s}_{0}-\lambda\overline{d}_{0}\geq\overline{s}_{k+1}-\lambda\overline{d}_{k+1}$. That completes the proof of the first statement.
However, when $\lambda(n)<\lambda\leq\lambda(n+1)$, from (31) and (32), we obtain that $\overline{s}_{n+1}-\lambda\overline{d}_{n+1}\geq\overline{s}_{k}-\lambda\overline{d}_{k}$, for all $k\geq 0$. Therefore, in this regime for $\lambda$, we can deduce that the optimal threshold value is $s_{n+1}$. That completes the proof of the second statement.
Let $\lim_{k\to\infty}\lambda(s_{k})=\lambda_{\infty}$. Since $\lambda(.)$ is increasing in $s_{k}$, then $\lambda(s_{k})<\lambda_{\infty}$ for all $k$. Thus, when $\lambda_{\infty}\leq\lambda$, from (30), we have $\overline{s}_{k}-\lambda\overline{d}_{k}<\overline{s}_{k+1}-\lambda\overline{d}_{k+1}$ as $\lambda(s_{k})<\lambda_{\infty}\leq\lambda$ for all $k$. Note that the optimal threshold $s_{n}$ cannot be finite, otherwise $\overline{s}_{n}-\lambda\overline{d}_{n}\geq\overline{s}_{n+1}-\lambda\overline{d}_{n+1}$, which is a contradiction. That completes the proof of the third statement. ∎
@@ III-A Problem Statement
Let $N$ be the total number of subsystems, $p_{i}$ be the probability that the packet sent by the source $i$ is successfully decoded, $r_{i}$ be the probability that the source $i$ transitions from one state to another, and $q_{i}$ be the probability that the adversary succeeds in jamming the channel $i$. Due to the energy constraint, the adversary can jam only $M<N$ channels. A jamming policy $\phi$ is defined as a sequence of actions $\phi=(\boldsymbol{d}_{\phi}(0),\boldsymbol{d}_{\phi}(1),\ldots)$, where $\boldsymbol{d}_{\phi}(t)=(d^{1}_{\phi}(t),d^{2}_{\phi}(t),\ldots,d^{N}_{\phi}(t))$ is a binary vector such that $d^{i}_{\phi}(t)=1$ if channel $i$ is jammed at time $t$. Let $\boldsymbol{s}(t)=(s^{1}(t),\ldots,s^{N}(t))$ be a vector at time $t$ such that $s^{i}(t)$ is the EAoII of subsystem $i$ at time $t$. Our aim is to find the optimal jamming policy that maximizes the total average EAoII subject to the constraint on the number of channels under attack. This can be formulated as
@@
$\text{subject to}\ \sum_{i=1}^{N}d^{i}_{\phi}(t)\leq\alpha N\quad\forall t\geq 0$ (33b)
@@ III-B Lagrangian Relaxation
subject $\text{subject to}\ \lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{\phi}\Big{[}\sum_{t=0}^{T-1}\sum_{i=1}^{N}d^{i}_{\phi}(t)\Big{]}\leq\alpha N.$ (34b)
@@
$f(W,\phi)$ $=\!\lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{\phi}\Big{[}\sum_{t=0}^{T-1}\sum_{i=1}^{N}s^{i}_{\phi}(t)\!-\!Wd^{i}_{\phi}(t)\big{|}\boldsymbol{s}(0)\Big{]}$
@@
$\underset{\phi\in\Phi}{\text{maximize}}\ \lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{\phi}\Big{[}\sum_{t=0}^{T-1}s^{i}_{\phi}(t)-Wd^{i}_{\phi}(t)\big{|}s^{i}(0)\Big{]}.$ (37)
@@ III-C Whittle’s Index Jamming Policy
and $d_{n}^{i}(W)$ be the action at state $s_{n}^{i}$ associated with subsystem $i$ under the optimal threshold jamming policy given the Lagrangian parameter $W$. We first consider the steady-state form of the problem in (37) under the threshold jamming policy with parameter $s^{i}_{n}$, i.e.,
$\underset{n\in\mathbb{N}}{\text{maximize}}\ \overline{s}_{n}^{i}-W\overline{d}_{n}^{i}$ (39)
where $\overline{s}_{n}^{i}$ and $\overline{d}_{n}^{i}$ are given in (2) and (26), respectively, when replacing $p$, $q$, and $r$ with $p_{i}$, $q_{i}$, $r_{i}$. To ensure the existence of the Whittle’s indices, it is imperative to establish the indexability property for all subsystems. To accomplish this, we initially formalize the concepts of indexability and the Whittle’s index in the following definitions.
@@
In the context of the problem in (37), for a given $W$, we define $D^{i}(W)$ as the set of states in which the optimal action for each subsystem $i$ is passive, i.e., $d^{i}_{n}(W)=0$. In other words, $s^{i}_{n}\in D^{i}(W)$ if and only if the optimal action at state $s^{i}_{n}$ is passive.
@@
To prove the claim, it is sufficient to show that $\overline{d}^{i}_{n}$ is decreasing with respect to $n$. Indeed, we can write
$\overline{d}^{i}_{n+1}-\overline{d}^{i}_{n}$
@@
Now that the indexability property has been established, we can assert the existence of the Whittle’s index. Building on the procedure introduced in [28, Algorithm 1], we propose a tailored algorithm that iteratively computes the Whittle’s index values for each subsystem $i$. The following lemma shows that the output of this algorithm is compatible with Definition 3.
@@
2:Find $W_{0}^{i}=\underset{n\in\mathbb{N}}{\text{inf}}\frac{\overline{s}_{n}^{i}-\overline{s}_{0}^{i}}{\overline{d}_{n}^{i}-\overline{d}_{0}^{i}}$
@@
8:     Find $W_{j}^{i}=\underset{n\in\mathbb{N}\setminus M_{j}^{i}}{\text{inf}}\frac{\overline{s}_{n}^{i}-\overline{s}_{n_{j}}^{i}}{\overline{d}_{n}^{i}-\overline{d}_{n_{j}}^{i}}$
@@
The proof follows directly the results in [28, Proposition 4]. ∎
In the next lemma, we provide a useful property that is satisfied by the sequences $\overline{s}_{k}^{i}$ and $\overline{d}_{k}^{i}$.
@@
For each subsystem $i$, if $\overline{d}_{k}^{i}$ is decreasing with $k$ and
$\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i}}<\frac{\overline{s}_{k+2}^{i}-\overline{s}_{k+1}^{i}}{\overline{d}_{k+2}^{i}-\overline{d}_{k+1}^{i}}$ (42)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{n}^{i}-\overline{d}_{k}^{i}}\geq\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i}}$ (43)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{n}^{i}-\overline{d}_{k}^{i}}=\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i}}\geq\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i}}$ (44)
@@
$\frac{\overline{s}_{n+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i}}=$ $\frac{\overline{s}_{n+1}^{i}-\overline{s}_{k}^{i}+\overline{s}_{n}^{i}-\overline{s}_{n}^{i}}{\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i}}$
$=$ $\frac{\overline{s}_{n+1}^{i}-\overline{s}_{n}^{i}}{\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i}}+\frac{\overline{s}_{n}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i}}$
$\geq$ $\underbrace{\frac{(\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i})(\overline{d}_{n+1}^{i}-\overline{d}_{n}^{i})}{(\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i})(\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i})}}_{(a)}+\underbrace{\frac{(\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i})(\overline{d}_{n}^{i}-\overline{d}_{k}^{i})}{(\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i})(\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i})}}_{(b)}$
$=$ $\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i}}\Big{[}\frac{\overline{d}_{n+1}^{i}-\overline{d}_{n}^{i}}{\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i}}+\frac{\overline{d}_{n}^{i}-\overline{d}_{k}^{i}}{\overline{d}_{n+1}^{i}-\overline{d}_{k}^{i}}\Big{]}$
$=$ $\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i}}$ (46)
where $(a)$ comes from (42) and the fact that $\overline{d}_{k}^{i}$ is strictly decreasing with respect to $k$, and (b) comes from (44) and the fact that $\overline{d}_{k}^{i}$ is decreasing with respect to $k$. Therefore, the inequality also holds for $n+1$. This concludes the proof. ∎
@@
$W(s_{n}^{i})=\lambda_{i}(s_{n})=\frac{\overline{s}_{n+1}^{i}-\overline{s}_{n}^{i}}{\overline{d}_{n+1}^{i}-\overline{d}_{n}^{i}}$ (47)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{j}^{i}}{\overline{d}_{n}^{i}-\overline{d}_{j}^{i}}\geq\frac{\overline{s}_{j+1}^{i}-\overline{s}_{j}^{i}}{\overline{d}_{j+1}^{i}-\overline{d}_{j}^{i}}$ (49)
@@
$\frac{\overline{s}_{k+1}^{i}-\overline{s}_{k}^{i}}{\overline{d}_{k+1}^{i}-\overline{d}_{k}^{i}}<\frac{\overline{s}_{k+2}^{i}-\overline{s}_{k+1}^{i}}{\overline{d}_{k+2}^{i}-\overline{d}_{k+1}^{i}}$ (50)
@@
$\frac{\overline{s}_{n}^{i}-\overline{s}_{j}^{i}}{\overline{d}_{n}^{i}-\overline{d}_{j}^{i}}\geq\frac{\overline{s}_{j+1}^{i}-\overline{s}_{j}^{i}}{\overline{d}_{j+1}^{i}-\overline{d}_{j}^{i}}$ (51)
for all $n>j$. Thus, the minimizer of $(\overline{s}_{n}^{i}-\overline{s}_{j}^{i})/(\overline{d}_{n}^{i}-\overline{d}_{j}^{i})$ over $n$ at step $j$ is $j+1$. As a result, the Whittle’s index at state $s_{j}^{i}$ is obtained as
$W(s_{j}^{i})=W_{j}^{i}=\frac{\overline{s}_{j+1}^{i}-\overline{s}_{j}^{i}}{\overline{d}_{j+1}^{i}-\overline{d}_{j}^{i}}=\lambda_{i}(s_{j}).$ (52)